OpenStudy (anonymous):

Factor completely: 2x3 + 6x2 + 2x + 6 2[(x + 3)(x2 + 1)] (2x + 6)(x2 + 1) (x + 3)(2x2 + 2) 2(x3 + 3x2 + x + 3)

OpenStudy (anonymous):

@dpasingh

OpenStudy (mathmate):

Note the pattern of the coefficients 2,6,2,6 which helps us factorize by grouping: \( 2x^3+6x^2+2x+6\\ =2(x^3+3x^2+x+3)\\ =2[x^2(x+3)+(x+3)] \) Can you continue with the factorization?

OpenStudy (anonymous):

Thank you lots.

OpenStudy (mathmate):

You're welcome!

OpenStudy (anonymous):

Actually that happens to not be an answer choice can it be wrote differently?

OpenStudy (anonymous):

@calculusxy

OpenStudy (anonymous):

\[2x^3 + 6x^2 + 2x + 6\]First factor out your 2 \[2 (x^3+3 x^2+x+3)\]\[2 (x^2+1)(x+3)\] (x^2+1) Got it from here?

OpenStudy (anonymous):

Dont mind that "(x^2+1)" at the bottom

OpenStudy (anonymous):

Thank I was confused about the first answer so how do I work out these kinds of problems?

OpenStudy (mathmate):

This is how you can work it out for similar problems: Note the pattern of the coefficients 2,6,2,6 which helps us factorize by grouping. First factor out the common factor 2 \(2x^3 +6x^2 +2x+6 =2(x^3 +3x^2\ \ +\ \ x+3) \) Now factor out x+3 in each of the two groups: \( 2[ x^2(x+3)+(x+3)] \) Notice that (x+3) is now a common factor, so we will factor it out to get \( 2x^2(x+3)(x^2+1) \) as the final answer.